Grade 12th passDiscuss with colleagues and IITians

Water is filled to a height H behind a dam of width w.The resultant force on dam is

Satya
8 Years agoGrade 12th pass

2 Answers

Arun
8 Years ago
Dear Satya
 
We know that pressure varies with height
So average pressure
P(avg) = (0 + d*g*H)/2 = dgH/2
Total force on the dam is-
F = P(avg)*A
F = (dgH/2) * (Hw)
F = dgH²w/2
d = density

Regards
Arun (askIITians forum expert)
Kushagra Madhukar
5 Years ago
Dear student,
Please find the solution to your problem.
 
We know that pressure varies with height
So average pressure
Pavg = (0 + ρgH)/2 = ρgH/2
Total force on the dam is-
F = Pavg x A
F = (ρgH/2) x (Hw)
F = ρgH2w/2
 
Thanks and regards,
Kushagra